# GeoBench: Benchmarking and Analyzing Monocular Geometry Estimation Models

### v1

@@ Abstract
Recent advances in discriminative and generative pretraining have yielded geometry estimation models with strong generalization capabilities. While discriminative monocular geometry estimation methods rely on large-scale fine-tuning data to achieve zero-shot generalization, several generative-based paradigms show the potential of achieving impressive generalization performance on unseen scenes by leveraging pre-trained diffusion models and fine-tuning on even a small scale of synthetic training data. Frustratingly, these models are trained with different recipes on different datasets, making it hard to find out the critical factors that determine the evaluation performance. Besides, the current widely used geometry evaluation benchmarks have two main drawbacks that may prevent the development of the field, i.e., limited scene diversity and unfavorable label quality. To resolve the above issues, (1) we build fair and strong baselines in a unified codebase for evaluating and analyzing the state-of-the-art (SOTA) geometry estimation models in terms of both different finetuning paradigms and training recipes; (2) we evaluate monocular geometry estimators on more challenging benchmarks for geometry estimation task with diverse scenes and high-quality annotations. Our results reveal that pre-trained using large data, discriminative models such as DINOv2, can outperform generative counterparts with a small amount of high-quality synthetic training data under the same training configuration, which suggests that fine-tuning data quality is a more important factor than the data scale and model architecture. Our observation also raises a question: if simply fine-tuning a general vision model such as DINOv2 using a small amount of synthetic depth data produces SOTA results, do we really need complex models, e.g., Marigold [32] and DepthFM [18] for depth estimation? We believe this work can propel advancements in geometry estimation tasks as well as a wide range of downstream applications.
@@ 1 Introduction
Although a large number of monocular geometry estimation models exist, they can be divided into two paradigms, i.e., discriminative-based and generative-based. Most contemporary discriminative monocular geometry estimation models leverage the pre-train priors from fully-supervised image classification backbones, e.g., ConvNeXt [63], EfficientNet [60] and ViT [7], or self-supervised backbones. e.g., DINOv2 [43], previous best deterministic depth estimation models scale-up their performance with pre-trained DINOv2 backbone and a large scale of fine-tuning data. Generative geometry estimation models [32, 17, 68, 18, 40] unleash the power of pre-trained text-to-image diffusion models, e.g., Stable Diffusion (SD) [46]. Several generative geometry estimation models [32, 17, 68, 18] show strong generation capability with even a small-scale high-quality synthetic fine-tuning data.
@@
To address the aforementioned problems, we perform a comprehensive geometry estimation benchmarking study from two perspectives. (1) Training strategy. We reimplement a bunch of SOTA algorithms in a unified codebase, including Marigold [32], Geowizard [17], GenPercept [68], DepthFM [18], DMP [40], Depth-Anything [69] and DSINE [1]. As such, we can fairly evaluate their performance under the same training configuration, and figure out whether the performance improvement is coming from the model architecture or coming from the high-quality training data. Previous generative geometry models are all based on Stable Diffusion 2.1 [46] with limited trained data (compared to the data scale of discriminative models), we further explore the potential of generative geometry models by conducting data scale-up and model scale-up ablations. (2) More benchmark datasets. Apart from traditional geometry evaluation benchmarks, we build more diverse scenes with high-quality labels for geometry evaluation. For depth estimation, we introduce three extra benchmark datasets, InSpaceType [65], MatrixCity [37], and Infinigen [51]. InSpaceType is an indoor depth evaluation benchmark, which contains 12 scenes, 1260 images, and resolution. It is a good complement for indoor benchmarks like NYUv2 and ScanNet. MatrixCity is a rendered dataset with real city-scale scenes, we select 808 street images and 403 aerial images for evaluation. It is suitable for evaluating driving and city scenes. Infinigen is also a high-quality rendered dataset, which contains diverse nature scenes. We use it to verify the generalization capability of depth estimation foundation models in wild scenes. For surface normal estimation, we expand existing benchmark datasets with more high quality and diverse datasets, e.g., indoor MuSHRoom dataset [53], outdoor Tank and Temples (T&T) dataset [33]111The surface normal annotation of MushRoom and T&T are obtained from Gaustudio [71], and wild Infinigen [51] dataset.
With the unified codebase, training data, and comprehensive benchmark datasets, we surprisingly find that (1) without bells and whistles, deterministic depth estimation model pre-trained on DINOv2 [43] backbone, followed by a simple DPT head [45], can already achieve great generalization performance by using only a small amount of high-quality synthetic datasets (K training samples). With the same training configuration, the deterministic model outperforms the generative counterpart, Marigold [32]. Besides, the result is contrary to previous deterministic models using large data to achieve generalization capability, e.g., Metric3Dv2 [24], which focuses on collecting more diverse training datasets (M training samples), and Depth-Anything [69], which focuses on scaling up model performance with large-scale pseudo labels (M training samples). Based on the result, we argue that training data quality is more important than data scale. (2) Generative-based geometry estimation models can generate high-resolution and detailed depth and surface normal maps, which is an advantage over the existing deterministic depth estimation models. We mainly attribute this to high-quality synthetic fine-tuning data and the design of the VAE [34] decoder, which can decode compressed depth latent to the same resolution depth map as the input image. (3) For surface normal estimation, the deterministic model DSINE [1] outperforms generative-based fine-tuning protocols with the same training configuration, which suggests apart from large-scale pre-training, inductive bias [1] is also an important factor in providing abundant information for surface normal estimation task. We hope our benchmarking results could pave the way for designing more powerful geometry estimation algorithms and developing more high-quality geometry estimation training datasets in the future.
@@ 2 Preliminaries
Deterministic geometry estimation models. With the widespread application of deep learning [35], learning-based methods have demonstrated their ability to estimate geometric information from monocular images [16, 20, 50, 69]. Early works primarily relied on deterministic models using either supervised or unsupervised methods. Eigen et al. [16] proposed the first learning-based method for monocular depth estimation, employing two deep network stacks and using ground truth depth for supervision. Zhou et al. proposed an early unsupervised framework, SfMLearner [75], in which camera pose and monocular depth are learned together. With the availability of large amounts of data, recent methods [50, 69, 74, 24] have shown a trend toward using large-scale datasets to develop robust geometry estimation models that generalize well to diverse environments. For instance, Ranftl et al. [50] introduced a method that demonstrates strong zero-shot testing ability by utilizing mixed training datasets. Yang et al. [69] further improved zero-shot monocular depth estimation performance by proposing Depth-Anything, which leverages large-scale data to achieve strong generalization ability. Meanwhile, Yin et al. [74, 24] proposed Metric3D, which can output accurate metric depth by training models on 11 public RGB-D datasets. Apart from depth estimation, advancements in surface normal information have also been achieved through the use of deterministic models. Surface normal information can not only be calculated directly from depth maps but can also be independently obtained through surface normal estimation techniques [64, 41, 38, 1]. For example, Bae et al. [1] proposed a method that demonstrates strong generalization capabilities and produces high-quality surface normal predictions by investigating inductive biases. Overall, the use of deterministic models for both depth and surface normal estimation has shown its significance in improving performance, thereby broadening the applications of monocular geometry estimation.
@@ 3.1 A Brief Overview of SOTA Methods
In a nutshell, discriminative models trained on large data, i.e., Depth-Anything [69], get the highest performance in most cases, while generative models finetuned on small data, e.g., Marigold [32], show competitive generalization capability on unseen scenes.
@@ 3.2 Benchmarking Different Generative Fine-Tuning Paradigms
Several fine-tuning paradigms have been proposed for diffusion-based depth estimation. Based on network architecture, they can be divided into two categories. The first category methods (Marigold [32] and DepthFM [18]) concatenate the image latent and depth latent encoded by VAE [34] encoder as the input of the UNet latent denoiser. As such, the input channels of the latent denoiser are doubled (8 input channels) to fit the expanded input. The second category methods (DMP [40] and GenPercept [68]) drop the depth latent, so they follow the original latent denoiser’s architecture (4 input channels). Based on fine-tuning paradigms, they can be divided into four categories. (1) Marigold [32] treats the initial depth latent as standard Gaussian noise and progressively denoise it with the same scheduler as the original Stable Diffusion pipeline. (2) DepthFM also treats the initial depth latent as standard Gaussian noise, however, the difference is that they finetune the denoiser with Flow Matching [36] pipeline, with auxiliary surface normal loss. (3) DMP [40] reformulates the task as a blending process, i.e., translating the image latent to depth latent with the Stable Diffusion v-prediction [46] learning target. (4) GenPercept [68] further improve the efficiency of DMP [40] by proposing a one-step inference pipeline. Based on the amount of fine-tuned parameters, they can be divided into two categories. The first category methods (Marigold, DepthFM, GenPercept) directly fine-tune the UNet parameters. The second category method (DMP) adds LORA [22] layers into the UNet architecture to achieve the goal of depth estimation.
@@ 3.3 Benchmarking the Inference Efficiency of Depth Estimation Foundation Models
Compared to deterministic models, the inference efficiency may become a bottleneck of the generative-based methods. In this section, we give detailed inference efficiency evaluation in Table 4. We can see that deterministic methods have fewer parameters than generative models. The main inference consumption of the generative models happens on VAE [34] and multiple inference steps of UNet. The last line of Table 4 shows that GenPercept [68] can achieve comparable inference latency with Depth-Anything (ViT-Large) and a tiny VAE encoder [42]. In Table 5, we found LCM [39] can effectively reduce the inference steps of Marigold [32] while maintaining the performance. Besides, a pre-trained tiny VAE [42] can substitute the standard VAE [46] with a minimal performance loss.
@@ 3.4 Benchmarking Discriminative and Generative Models on the Same Training Data
Can discriminative depth estimation models achieve competitive results with small-scale high-quality training datasets like generative-based methods? To answer this question, we benchmark deterministic and generative geometry model with the same amount of training data and the same training strategy. Specifically, we use three training datasets, i.e., Hypersim (38,387) [52], Virtual Kitti (16,790) [21] and Tartanair (31,008) [67], with total 77,897 samples. Both models are trained with 20,000 iterations, with a total batch size of 96 on 4 GPUs. For the deterministic depth model, we follow the network architecture of Depth-Anything [69] (ViT-Large backbone pre-trained with DINOv2 and DPT [45] head), supervised with the affine-invariant loss [69]. For the generative geometry model, we choose Marigold [32] as our baseline. We can see from Table 6 that (1) the discriminative model is largely inferior to generative-based Marigold on all evaluation datasets without DINOv2 pre-train (line 1 v.s.line 3). However, the discriminative model beats Marigold by a large margin when initialized with DINOv2 pre-train weight (line 2 v.s.line 3); (2) scale-up Marigold from SD21 to SDXL brings consistent improvement in all benchmarks. We can see from Table 7 that our deterministic model trained on 77K data outperforms Metric3Dv2 [24] in all three datasets, and, is comparable with Depth-Anything [69] in two datasets (InspaceType and Infinigen). This phenomenon suggests that high-quality fine-tuning data, rather than large-scale training data, is indispensable for discriminative models to achieve strong generalizable performance. The qualitative visualization of deterministic models is not as good as generative models, more visualizations are available in the Supp. Mat.
@@ 4.1 A Brief Overview of SOTA Methods
DSINE [1] and Metric3Dv2 [24] are two representative discriminative surface estimation models, which leverage the geometry priors from two distinct perspectives. DSINE leverages two forms of inductive bias: (1) per-pixel ray direction, and (2) the relationship between the neighboring surface normal, to learn a generalizable surface normal estimator. Metric3Dv2 [24] proposes to optimize the surface normal map by distilling diverse data knowledge from the estimated metric depth. Different from deterministic models, GeoWizard [17] is a generative surface normal estimator without using any inductive bias from the geometry priors. It purely relies on pre-trained diffusion priors to estimate the surface normal map. Table 9 summarizes their performance on six benchmarks. The Mushroom [53] (indoor), T&T [33] (outdoor), and Infinigen [51] (wild) datasets are constructed by us to add more diverse scenes with accurate surface normal labels in the evaluation benchmarks. We can see that Metric3Dv2 [24] outperform DSINE [1] and GeoWizard [17] in most datasets. Note it is an unfair comparison since (1) Metric3Dv2 [24] is trained on 16M images, while DSINE is trained on 160K images, and GeoWizard is trained on 280K images. (2) DSINE use a much smaller backbone, EfficientNet-B5 [60], while Metric3Dv2 [24] employs the ViT-Large [7] backbone, pretrained using DINOv2 with registers [12].
@@ 4.2 Benchmarking Discriminative and Generative Models on the Same Training Data
In this section, we fairly benchmark deterministic DSINE [1] and several representative generative geometry models, i.e., Marigold [32], DMP [40], GenPercept [68], and DepthFM [18], with 5 training datasets, Hypersim [52] (), Tartanair [67] (), Virtual Kitti [21] (), BlendedMVS [70] (), ClearGrasp [56] (), a total of samples. For generatative-based models, we represent the output surface normals as unit vectors. We follow DSINE [1] to represent the outputs of deterministic-based model as axis-angles with three degrees of freedom. All models are trained with iterations, batch sizes, resolution on 4 A800 GPUs. All generative-based models use learning rate. For deterministic model, we follow DSINE [1] to use learning rate for the backbone and learning rate for the decoder.
We can see from Table 8 that (1) DSINE can scale up the performance by using ViT-Large backbone with DINOv2 pretrain (compared with ImageNet pretrained Efficient-B5 backbone). (2) For generative-based fine-tuning protocols, DepthFM [18] outperforms other paradigms in most benchmarks. We attribute this to the decoder supervision during the training. Paradigms that requires multi-step denoising inference steps, e.g., Marigold [32] and DMP [40]are not suitable to perform decoder supervision during the training. (3) Deterministic models, equipped with inductive bias, outperform generative-based models with only diffusion priors. It is promising to inject inductive bias into the diffusion-based models, as such, the surface normal estimator can effectively leverage the diffusion priors and inductive bias to boost the performance. (4) DSINE (ViT-Large in Table 8) trained with K samples achieves comparable performance with Metric3Dv2 trained with M samples (Table 9). The results verify the point that data-quality is more important than the data-scale.

### v2

@@ Abstract
Recent advances in discriminative and generative pretraining have yielded geometry estimation models with strong generalization capabilities. While discriminative monocular geometry estimation methods rely on large-scale fine-tuning data to achieve zero-shot generalization, several generative-based paradigms show the potential of achieving impressive generalization performance on unseen scenes by leveraging pre-trained diffusion models and fine-tuning on even a small scale of synthetic training data. Frustratingly, these models are trained with different recipes on different datasets, making it hard to find out the critical factors that determine the evaluation performance. Besides, the current widely used geometry evaluation benchmarks have two main drawbacks that may prevent the development of the field, i.e., limited scene diversity and unfavorable label quality. To resolve the above issues, (1) we build fair and strong baselines in a unified codebase for evaluating and analyzing the state-of-the-art (SOTA) geometry estimation models in terms of both different finetuning paradigms and training recipes; (2) we evaluate monocular geometry estimators on more challenging benchmarks for geometry estimation task with diverse scenes and high-quality annotations.
Our results reveal that pre-trained using large data, discriminative models such as DINOv2, can outperform generative counterparts with a small amount of high-quality synthetic training data under the same training configuration, which suggests that fine-tuning data quality is a more important factor than the data scale and model architecture. Our observation also raises a question: if simply fine-tuning a general vision model such as DINOv2 using a small amount of synthetic depth data produces SOTA results, do we really need complex models, e.g., Marigold [32] and DepthFM [18] for depth estimation? We believe that this work can propel advancements in geometry estimation tasks and a wide range of other downstream vision tasks.
@@ 1 Introduction
Although a large number of monocular geometry estimation models exist, they can be divided into two paradigms, i.e., discriminative-based and generative-based. Most contemporary discriminative monocular geometry estimation models leverage the pre-train priors from fully-supervised image classification backbones, e.g., ConvNeXt [63], EfficientNet [60] and ViT [7], or self-supervised backbones. e.g., DINOv2 [43], previous best discriminative depth estimation models scale-up their performance with pre-trained DINOv2 backbone and a large scale of fine-tuning data. Generative geometry estimation models [32, 17, 68, 18, 40] unleash the power of pre-trained text-to-image diffusion models, e.g., Stable Diffusion (SD) [46]. Several generative geometry estimation models [32, 17, 68, 18] show strong generation capability with even a small-scale high-quality synthetic fine-tuning data.
@@
To address the aforementioned problems, we perform a comprehensive geometry estimation benchmarking study from two perspectives:
1.
@@
Training strategy. We reimplement a comprehensive list of SOTA algorithms in a unified codebase, including Marigold [32], Geowizard [17], GenPercept [68], DepthFM [18], DMP [40], Depth-Anything [69] and DSINE [1]. As such, we can fairly evaluate their performance under the same training configuration, and figure out whether the performance improvement is coming from the model architecture or coming from the high-quality training data. Previous generative geometry models are all based on Stable Diffusion 2.1 [46] with limited trained data (compared to the data scale of discriminative models), we further explore the potential of generative geometry models by conducting data scale-up and model scale-up ablations.
2.
More benchmark datasets. Apart from traditional geometry evaluation benchmarks, we build more diverse scenes with high-quality labels for geometry evaluation. For depth estimation, we introduce three extra benchmark datasets, InSpaceType [65], MatrixCity [37], and Infinigen [51]. InSpaceType is an indoor depth evaluation benchmark, which contains 12 scenes, 1260 images, and resolution. It is a good complement for indoor benchmarks like NYUv2 and ScanNet. MatrixCity is a rendered dataset with real city-scale scenes, we select 808 street images and 403 aerial images for evaluation. It is suitable for evaluating driving and city scenes. Infinigen is also a high-quality rendered dataset, which contains diverse nature scenes. We use it to verify the generalization capability of depth estimation foundation models in wild scenes. For surface normal estimation, we expand existing benchmark datasets with more high quality and diverse datasets, e.g., indoor MuSHRoom dataset [53], outdoor Tank and Temples (T&T) dataset [33]111The surface normal annotation of MushRoom and T&T are obtained from Gaustudio [71], and wild Infinigen [51] dataset.
With the unified codebase, training data, and comprehensive benchmark datasets, we surprisingly find that:
1.
Without bells and whistles, the depth estimation model pre-trained on the DINOv2 [43] backbone, followed by a simple DPT head [45], can already achieve great generalization performance by using only a small amount of high-quality synthetic datasets (K training samples).
With the same training configuration, the discriminative model outperforms the recent generative counterpart, Marigold [32]. Besides, the result is a sharp contrast to previous discriminative models using large data to achieve generalization capability, e.g., Metric3Dv2 [24], which focuses on collecting more diverse training datasets (M training samples), and Depth-Anything [69], which focuses on scaling up model performance with large-scale pseudo labels (M training samples). Based on the result, we argue that training data quality may be more important than the data scale.
2.
Generative-based geometry estimation models can generate high-resolution and detailed depth and surface normal maps, which is an advantage over the existing discriminative depth estimation models. We mainly attribute this to high-quality synthetic fine-tuning data and the design of the VAE [34] decoder, which can decode compressed depth latent to the same resolution depth map as the input image.
3.
For surface normal estimation, the discriminative model DSINE [1] outperforms generative-based fine-tuning protocols with the same training configuration, suggesting that, apart from large-scale pre-training, inductive bias [1] can also be an important factor in providing rich information for surface normal estimation tasks.
We hope that our benchmarking results can pave the way for designing more powerful geometry estimation algorithms and developing more high-quality geometry estimation training datasets in the future.
@@ 2 Preliminaries
Discriminative geometry estimation models. With the widespread application of deep learning [35], learning-based methods have demonstrated their ability to estimate geometric information from monocular images [16, 20, 50, 69]. Early works primarily relied on discriminative models using either supervised or unsupervised methods. Eigen et al. [16] proposed the first learning-based method for monocular depth estimation, employing two deep network stacks and using ground truth depth for supervision. Zhou et al. proposed an early unsupervised framework, SfMLearner [75], in which camera pose and monocular depth are learned together. With the availability of large amounts of data, recent methods [50, 69, 74, 24] have shown a trend toward using large-scale datasets to develop robust geometry estimation models that generalize well to diverse environments. For instance, Ranftl et al. [50] introduced a method that demonstrates strong zero-shot testing ability by utilizing mixed training datasets. Yang et al. [69] further improved zero-shot monocular depth estimation performance by proposing Depth-Anything, which leverages large-scale data to achieve strong generalization ability. Meanwhile, Yin et al. [74, 24] proposed Metric3D, which can output accurate metric depth by training models on 11 public RGB-D datasets. Apart from depth estimation, advancements in surface normal information have also been achieved through the use of discriminative models. Surface normal information can not only be calculated directly from depth maps but can also be independently obtained through surface normal estimation techniques [64, 41, 38, 1]. For example, Bae et al. [1] proposed a method that demonstrates strong generalization capabilities and produces high-quality surface normal predictions by investigating inductive biases. Overall, the use of discriminative models for both depth and surface normal estimation has shown its significance in improving performance, thereby broadening the applications of monocular geometry estimation.
@@ 3.1 A Brief Overview of SOTA Methods
In a nutshell, discriminative models trained on large data, i.e., Depth-Anything [69], obtains the highest performance in most cases, while generative models fine-tuned on small data, e.g., Marigold [32], show competitive generalization capability on unseen scenes.
@@ 3.2 Benchmarking Different Generative Fine-Tuning Paradigms
Several fine-tuning paradigms have been proposed for diffusion-based depth estimation. Based on the network architecture, they can be divided into two categories. The first category of methods (Marigold [32] and DepthFM [18]) concatenate the image latent and depth latent encoded by VAE [34] encoder as the input of the UNet latent denoiser. As such, the input channels of the latent denoiser are doubled (8 input channels) to fit the expanded input. The second category methods (DMP [40] and GenPercept [68]) drop the depth latent, so they follow the original latent denoiser’s architecture (4 input channels). Based on fine-tuning paradigms, they can be divided into four categories. (1) Marigold [32] treats the initial depth latent as standard Gaussian noise and progressively denoise it with the same scheduler as the original Stable Diffusion pipeline. (2) DepthFM also treats the initial depth latent as standard Gaussian noise, however, the difference is that they finetune the denoiser with Flow Matching [36] pipeline, with auxiliary surface normal loss. (3) DMP [40] reformulates the task as a blending process, i.e., translating the image latent to depth latent with the Stable Diffusion v-prediction [46] learning target. (4) GenPercept [68] further improve the efficiency of DMP [40] by proposing a one-step inference pipeline. Based on the amount of fine-tuned parameters, they can be divided into two categories. The first category methods (Marigold, DepthFM, GenPercept) directly fine-tune the UNet parameters. The second category method (DMP) adds LORA [22] layers into the UNet architecture to achieve the goal of depth estimation.
@@ 3.3 Benchmarking the Inference Efficiency of Depth Estimation Foundation Models
Compared to discriminative models, the inference efficiency may become a bottleneck of the generative-based methods. In this section, we give detailed inference efficiency evaluation in Table 4. We can see that discriminative methods have fewer parameters than generative models. The main inference consumption of the generative models happens on VAE [34] and multiple inference steps of UNet. The last line of Table 4 shows that GenPercept [68] can achieve comparable inference latency with Depth-Anything (ViT-Large) and a tiny VAE encoder [42]. In Table 5, we found LCM [39] can effectively reduce the inference steps of Marigold [32] while maintaining the performance. Besides, a pre-trained tiny VAE [42] can substitute the standard VAE [46] with a minimal performance loss.
@@ 3.4 Benchmarking Discriminative and Generative Models on the Same Training Data
Can discriminative depth estimation models achieve competitive results with small-scale high-quality training datasets like generative-based methods? To answer this question, we benchmark discriminative and generative geometry model with the same amount of training data and the same training strategy. Specifically, we use three training datasets, i.e., Hypersim (38,387) [52], Virtual Kitti (16,790) [21] and Tartanair (31,008) [67], with total 77,897 samples. Both models are trained with 20,000 iterations, with a total batch size of 96 on 4 GPUs. For the discriminative depth model, we follow the network architecture of Depth-Anything [69] (ViT-Large backbone pre-trained with DINOv2 and DPT [45] head), supervised with the affine-invariant loss [69]. For the generative geometry model, we choose Marigold [32] as our baseline.
We can see from Table 6 that (1) the discriminative model is largely inferior to generative-based Marigold on all evaluation datasets without DINOv2 pre-train (line 1 vs. line 3). However, the discriminative model beats Marigold by a large margin when initialized with DINOv2 pre-train weight (line 2 vs. line 3); (2) scale-up Marigold from SD21 to SDXL brings consistent improvement in all benchmarks. We can see from Table 7 that our discriminative model trained on 77K data outperforms Metric3Dv2 [24] in all three datasets, and, is comparable with Depth-Anything [69] on two datasets (InspaceType and Infinigen). This phenomenon suggests that high-quality fine-tuning data, rather than large-scale training data, is indispensable for discriminative models to achieve strong generalizable performance. The qualitative visualization of discriminative models is not as good as generative models, more visualizations are available in the Supp. Mat.
@@ 4.1 A Brief Overview of SOTA Methods
DSINE [1] and Metric3Dv2 [24] are two representative discriminative surface estimation models, which leverage the geometry priors from two distinct perspectives. DSINE leverages two forms of inductive bias: (1) per-pixel ray direction, and (2) the relationship between the neighboring surface normal, to learn a generalizable surface normal estimator. Metric3Dv2 [24] proposes to optimize the surface normal map by distilling diverse data knowledge from the estimated metric depth. Different from discriminative models, GeoWizard [17] is a generative surface normal estimator without using any inductive bias from the geometry priors. It purely relies on pre-trained diffusion priors to estimate the surface normal map. Table 9 summarizes their performance on six benchmarks. The Mushroom [53] (indoor), T&T [33] (outdoor), and Infinigen [51] (wild) datasets are constructed by us to add more diverse scenes with accurate surface normal labels in the evaluation benchmarks. We can see that Metric3Dv2 [24] outperform DSINE [1] and GeoWizard [17] in most datasets. Note it is an unfair comparison since (1) Metric3Dv2 [24] is trained on 16M images, while DSINE is trained on 160K images, and GeoWizard is trained on 280K images. (2) DSINE use a much smaller backbone, EfficientNet-B5 [60], while Metric3Dv2 [24] employs the ViT-Large [7] backbone, pretrained using DINOv2 with registers [12].
@@ 4.2 Benchmarking Discriminative and Generative Models on the Same Training Data
In this section, we fairly benchmark discriminative DSINE [1] and several representative generative geometry models, i.e., Marigold [32], DMP [40], GenPercept [68], and DepthFM [18], with 5 training datasets, Hypersim [52] (), Tartanair [67] (), Virtual Kitti [21] (), BlendedMVS [70] (), ClearGrasp [56] (), a total of samples. For generatative-based models, we represent the output surface normals as unit vectors. We follow DSINE [1] to represent the outputs of discriminative-based model as axis-angles with three degrees of freedom. All models are trained with iterations, batch sizes, resolution on 4 A800 GPUs. All generative-based models use learning rate. For discriminative model, we follow DSINE [1] to use learning rate for the backbone and learning rate for the decoder.
We can see from Table 8 that (1) DSINE can scale up the performance by using ViT-Large backbone with DINOv2 pretrain (compared with ImageNet pretrained Efficient-B5 backbone). (2) For generative-based fine-tuning protocols, DepthFM [18] outperforms other paradigms in most benchmarks. We attribute this to the decoder supervision during the training. Paradigms that requires multi-step denoising inference steps, e.g., Marigold [32] and DMP [40]are not suitable to perform decoder supervision during the training. (3) Discriminative models, equipped with inductive bias, outperform generative-based models with only diffusion priors. It is promising to inject inductive bias into the diffusion-based models, as such, the surface normal estimator can effectively leverage the diffusion priors and inductive bias to boost the performance. (4) DSINE (ViT-Large in Table 8) trained with K samples achieves comparable performance with Metric3Dv2 trained with M samples (Table 9). The results verify the point that data-quality is more important than the data-scale.